### v1

# Open cell property in weakly o-minimal structures

Abstract.
In an o-minimal structure, Wilkie proved that every bounded definable open set is a finite union of definable open cells. We generalze it to weakly o-minimal structures.
Key words and phrases:
Weakly o-minimal structures, open strong cells and open cell property
2020 Mathematics Subject Classification:
Primary 03C64
1. Introduction
Let $\mathcal{M}=(M,<,\ldots)$ be an expansion of a dense linear order without endpoints. A subset $A$ of $M$ is said to be convex if $a,b\in A$ and $c\in M$ with $a<c<b$ then $c\in A$. Moreover if $A=\emptyset$ or $\inf A,{\rm sup}A\in M\cup\{-\infty,+\infty\}$, then $A$ is called an interval in $M$. We say that $\mathcal{M}$ is o-minimal (weakly o-minimal) if every definable subset of $M$ is a finite union of intervals (convex sets), respectively.
In o-minimal structures, Wilkie ([9]) proved every bounded definable open set is a finite union of open cells. Andrews ([1]) proved that if an o-minimal structure admits CE-cell decomposition then any definable open set is expressed as a finite union of definable open cells. Edmundo et al ([3]) proved in semi-bounded o-minimal expansion of an ordered group, any definable open set is covered a finite union of open cells. The relation between strong cell decomposition in weakly o-minimal structures and open cell property in o-minimal structures is proved by [4]. We consider a generalization of the above results in weakly o-minimal structures.
2. Weakly o-minimal expansions of real closed fields
Let $\mathcal{M}=(M,<,\ldots)$ be an weakly o-minimal structure. For any subsets $C,D$ of $M$, we write $C<D$ if $c<d$ whenever $c\in C$ and $d\in D$. A pair $\langle C,D\rangle$ of non-empty subsets of $M$ is called a cut in $\mathcal{M}$ if $C<D,C\cup D=M$ and $D$ has no lowest element. A cut $\langle C,D\rangle$ is said to be definable in $\mathcal{M}$ if the sets $C,D$ are definable in $\mathcal{M}$. The set of all cuts definable in $\mathcal{M}$ will be denoted by $\overline{M}$. Note that we have $M=\overline{M}$ if $\mathcal{M}$ is o-minimal. We define a linear ordering on $\overline{M}$ by $\langle C_{1},D_{1}\rangle<\langle C_{2},D_{2}\rangle$ if and only if $C_{1}\subsetneq C_{2}$. Then we may treat $(M,<)$ as a substructure of $(\overline{M},<)$ by identifying an element $a\in M$ with the definable cut $\langle(-\infty,a],(a,+\infty)\rangle$. Morever, if $\mathcal{M}=(M,<,+,\ldots)$ is an expansion of an ordered group $(M,<,+)$, then the cut $\langle C,D\rangle$ is said to be valuational if there exists $\varepsilon>0$ such that for all $x\in C$ and $y\in D$, we have $y-x>\varepsilon$. Otherwise the cut $\langle C,D\rangle$ is said to be non-valuational. The structure $\mathcal{M}$ is said to be non-valuational if all definable cuts in $\mathcal{M}$ are non-valuational.
We equip $M$ ($\overline{M}$) with the interval topology (the open intervals form a base), and each product $M^{n}$ ($(\overline{M})^{n}$) with the corresponding product topology, respectively.
Recall the notion of definable functions from [8]. Let $n$ be a positive integer and $A\subseteq M^{n}$ definable. A function $f:A\to\overline{M}$ is said to be definable if the set $\{\langle x,y\rangle\in M^{n+1}:x\in A,y<f(x)\}$ is definable. A function $f:A\to\overline{M}\cup\{-\infty,+\infty\}$ is said to be definable if $f$ is a definable function from $A$ to $\overline{M}$, $f(x)=$-$\infty$ for all $x\in A$, or $f(x)=+\infty$ for all $x\in A$. A definable function $f:A\to\overline{M}\cup\{-\infty,+\infty\}$ is said to be regular if it takes all its values in one of the sets $M$, $\overline{M}-M$, $\{-\infty\}$, or $\{+\infty\}$.
Definition 2.1 ([7]).
For every $m\in\mathbb{N}_{+}$, we define, by induction, strong cells $C\subseteq M^{m}$ and their completions $\overline{C}\subseteq\overline{M}^{m}$.
  1. (1)
    Any singleton of $M$ is a 0-strong cell in $M$ and is equal to its completion.
  2. (2)
    Any non-empty open convex definable subset of $M$ is a 1-strong cell in $M$. If $C\subseteq M$ is a 1-strong cell, then we define its completion by $\overline{C}:=\{x\in\overline{M}:(\exists a,b\in C)(a<x<b)\}$.
    Let $m\in\mathbb{N}_{+}$, $k\leq m$ and suppose that we have already defined $k$-strong cells in $M^{m}$ and their completions in $\overline{M}^{m}$.
  3. (3)
    If $C\subseteq M^{m}$ is a $k$-strong cell and $f:C\to M$ is a continuous definable function which has a unique continuous extension $\overline{f}:\overline{C}\to\overline{M}$, then the graph of $f$ which is denoted by $\Gamma(f)_{C}$, is a $k$-strong cell in $M^{m+1}$ and its completion in $\overline{M}^{m+1}$ is defined as $\Gamma(\overline{f})_{\overline{C}}$.
  4. (4)
    If $C\subseteq M^{m}$ is a $k$-strong cell and $f,g:C\to M\cup\{\pm\infty\}$ is continuous definable functions such that $f,g$ have continuous extensions $\overline{f},\overline{g}:\overline{C}\to\overline{M}$, where $\overline{f}(\overline{a})<\overline{g}(\overline{a})$ for any $\overline{a}\in\overline{C}$, then the set $(f,g)_{C}:=\{\langle\overline{a},b\rangle\in C\times M:f(\overline{a})<b<g(\overline{a})\}$ is called a $(k+1)$-strong cell in $M^{m+1}$. The completion of $(f,g)_{C}$ in $\overline{M}^{m+1}$ is defined as $(\overline{f},\overline{g})_{\overline{C}}:=\{\langle\overline{a},b\rangle\in\overline{C}\times\overline{M}:\overline{f}(\overline{a})<b<\overline{g}(\overline{a})\}$.
A strong cell $C\subseteq M^{m}$, $m\geq 2$, is called a refined strong cell if each of the boundary functions appearing in its definition is regular. Refined strong cells in $M$ coincide with strong cells in $M$.
Let $C\subseteq M^{m}$ be a strong cell and $f:C\to\overline{M}$ definable. The function $f$ is said to be strongly continuous if it has a unique continuous extension $\overline{f}:\overline{C}\to\overline{M}$.
Definition 2.2.
Any finite partition of $M$ into singletons and convex open sets definable in $\mathcal{M}$ is called a strong cell decomposition of $M$. A finite partition $\mathcal{C}$ of $M^{m+1}$ into strong cells is said to be a strong cell decomposition of $M^{m+1}$ if $\pi[\mathcal{C}]:=\{\pi[C]:C\in\mathcal{C}\}$ is a strong cell decomposition of $M^{m}$. Here $\pi:M^{m+1}\to M^{m}$denotes the projection dropping the last coordinate. A strong cell decomposition $\mathcal{C}$ of $M^{m}$ partitions $X\subseteq M^{m}$ if for any $C\in\mathcal{C}$, we have $C\subseteq X$ or $C\cap X=\emptyset$. The structure $\mathcal{M}$ has the strong cell decomposition property if for any positive integers $m$, $k$ and any definable sets $X_{1},\ldots,X_{k}\subseteq M^{m}$, there exists a decomposition of $M^{m}$ into finitely many strong cells partitioning each of the sets $X_{1},\ldots,X_{k}$. Similarly, we can define the refined strong cell decomposition property.
The following result follows from Lemma 2.6 and Corollary 2.16 of [7].
Fact 2.3.
Let $\mathcal{M}=(M,<,+,\ldots)$ be a weakly o-minimal expansion of an ordered group. Then the structure $\mathcal{M}$ is a non-valuational if and only if it has a strong cell decomposition property.
Let $\mathcal{C}$ and $\mathcal{D}$ are refined strong cell decompositions of $M^{m}$. We say that $\mathcal{D}$ refines $\mathcal{C}$ (denoted by $\mathcal{C}\prec\mathcal{D}$) if every cell in $\mathcal{C}$ is a union of some cells in $\mathcal{D}$
The following statement is Fact 2.1 of [8].
Fact 2.4.
Let $k,m\in\mathbb{N}_{+}$. If $X_{1},\ldots,X_{k}\subseteq M^{m}$ are definable sets, then there exists a smallest refined strong cell decomposition $\mathcal{D}$ of $M^{m}$ partitioning each of $X_{1},\ldots,X_{k}$.
For any $n\in\mathbb{N}_{+}$ and any refined strong cell $C\subseteq M^{n}$, we denote by $R_{C}$ an $n$-ary relational symbol. We interpret $R_{C}$ in $\overline{M}^{n}$ as $\overline{C}$, the completion of $C$. According to Section 2 of [8], the structure $\overline{\mathcal{M}}:=(\overline{M},<,(R_{C}:C\text{ is a refined strong cell}))$ is o-minimal and is called the canonical o-minimal extension of $\mathcal{M}$. If $X\subseteq\overline{M}^{m}$ is a set definable in $\overline{\mathcal{M}}$, then $X\cap M^{m}$ is definable in $\mathcal{M}$. If additionally $Y\subseteq M^{m}$ is definable in $\mathcal{M}$, then $X\cap Y$ is definable in $\mathcal{M}$.
The following definition is Definition 3.1 of [6].
Definition 2.5.
Let $m\in\mathbb{N}_{+}$. Suppose that $X$ is a defianble subset of $M^{m}$ and $\mathcal{C}$ is the smallest refined strong cell decomposition of $M^{m}$ partitioning $X$. We define the completion of $X$ in $\overline{M}^{m}$ as $\overline{X}:=\bigcup\{\overline{C}:C\in\mathcal{C}\wedge C\subseteq X\}$.
Remark 2.6.
Note that there exist sets $C,D$ such that $C\subseteq D$ but $\overline{C}\not\subseteq\overline{D}$. Let $R_{alg}$ denote the ordered field of real algebraic numbers and let the unary predicate symbol $P$ be interpreted by the convex set $S=(-\pi,\pi)\cap R_{alg}$. Then, by Proposition 2.1 from [5], the structure $\mathfrak{R}=(R_{alg},<,+,\cdot,P)$ is non-valuational weakly o-minimal. So the structure $\mathfrak{R}$ has the refined strong cell decomposition property. Consider the functions $f,g:I=(0,4)\to\mathfrak{R}$ defined by $f(x)=x$ and $g(x)=x+1$, and the function $h:J=(0,\pi)\to\mathfrak{R}$ defined by $h(x)=x+2$. Let $C=\Gamma(f)_{I}\cup\Gamma(g)_{I}$ and $D=\Gamma(f)_{I}\cup\Gamma(g)_{I}\cup\Gamma(h)_{J}$. Then, we have $C\subseteq D$ but $\overline{C}\not\subseteq\overline{D}$.
Remark 2.7.
Let $m\in\mathbb{N}_{+}$. Suppose that $X$ is a defianble subset of $M^{m}$. Let $\mathcal{C}$ be the smallest refined strong cell decomposition of $M^{m}$ partitioning $X$ such that $X=\bigcup_{i=1}^{k}C_{i}$ and $\overline{X}=\bigcup_{i=1}^{k}\overline{C_{i}}$ with $C_{1},\ldots,C_{k}\in\mathcal{C}$. If $X$ is open in $M$, then $\overline{X}$ is open in $\overline{M}$.
Proposition 2.8 (Weak definable choice).
Suppose that $\mathcal{M}=(M,<,+,\ldots)$ is a weakly o-minimal expansion of an ordered group $(M,<,+)$. Let $S\subseteq M^{m+n}$ be definable and $\pi:M^{m+n}\to M^{m}$ the projection of the first $m$ coordinates. Then, there exists a definable map $f:\pi(S)\to\overline{M}^{n}$ such that $\Gamma(f)\subseteq\overline{S}$.
Proof.
Let $\mathcal{C}$ be the smallest refined strong cell decomposition of $M^{m+n}$ partitioning $S$ such that $S=\bigcup_{i=1}^{k}C_{i}$ and $\overline{S}=\bigcup_{i=1}^{k}\overline{C_{i}}$ with $C_{1},\ldots,C_{k}\in\mathcal{C}$. By the o-minimality of $\overline{{\mathcal{M}}}$ and Proposition 6.1.2 of [2], for any $C\in\mathcal{C}$ with $C\subseteq S$, there exists a map $\overline{f}_{C}:\overline{\pi(C)}\to\overline{M}^{n}$ such that $\Gamma(\overline{f}_{C})\subseteq\overline{C}$ and $\overline{f}_{C}$ is definable in $\overline{{\mathcal{M}}}$. By Propisition 2.3 from [8], $\overline{f}_{C}|_{\overline{\pi(C)}\cap M^{m}}:\overline{\pi(C)}\cap R^{m}\to\overline{M}^{n}$ is definable in $\mathcal{M}$}. By $S=\bigcup_{i=1}^{k}C_{i}$ and $\overline{S}=\bigcup_{i=1}^{k}\overline{C_{i}}$, this completes the proof. ∎
Proposition 2.9.
Suppose that $\mathcal{M}=(M,<,+,\ldots)$ is a weakly o-minimal expansion of an ordered group $(M,<,+)$. Let $S\subseteq M^{m+n}$ be definable and $\overline{\pi}:\overline{M}^{m+n}\to\overline{M}^{m}$ the projection of the first $m$ coordinates. Then, there exists a map $\overline{g}:\overline{\pi}(\overline{S})\to\overline{S}$ such that $\overline{\pi}\circ\overline{g}=\mathrm{id}_{\overline{\pi}(\overline{S})}$ and $\overline{g}$ is definable in $\overline{{\mathcal{M}}}$.
Proof.
Let $\pi:=\overline{\pi}|_{M^{m+n}}$. Suppose that $\mathcal{C}$ is the smallest refined strong cell decomposition of $M^{m+n}$ partitioning $S$ such that $S=\bigcup_{i=1}^{k}C_{i}$ and $\overline{S}=\bigcup_{i=1}^{k}\overline{C_{i}}$ with $C_{1},\ldots,C_{k}\in\mathcal{C}$. By the o-minimality of $\overline{{\mathcal{M}}}$ and Proposition 6.1.2 of [2], for any $C\in\mathcal{C}$ with $C\subseteq S$, there exists a map $\overline{f}_{C}:\overline{\pi(C)}\to\overline{M}^{n}$ such that $\Gamma(\overline{f}_{C})\subseteq\overline{C}$ and $\overline{f}_{C}$ is definable in $\overline{{\mathcal{M}}}$. Let $\overline{f}=\bigcup_{i=1}^{k}\overline{f}_{C_{i}}$. Then we have $\overline{f}:\overline{\pi}(\overline{S})\to\overline{M}^{n}$ and $\Gamma(\overline{f})\subseteq\overline{S}$. We define the $\overline{\mathcal{M}}$-definable map $\overline{g}:\overline{\pi}(\overline{S})\to\overline{S}$ by $\overline{g}(x)=(x,\overline{f}(x))$. Then it follows $\overline{\pi}\circ\overline{g}=\mathrm{id}_{\overline{\pi}(\overline{S})}$. ∎
Example 2.10.
Let $R_{alg}$ denote the ordered field of real algebraic numbers and let the unary predicate symbol $P$ be interpreted by the convex set $S=(-\pi,\pi)\cap R_{alg}$. Then, by Proposition 2.1 from [5], the structure $\mathfrak{R}=(R_{alg},<,+,\cdot,P)$ is non-valuational weakly o-minimal. So the structure $\mathfrak{R}$ has the refined strong cell decomposition property. Consider the functions $f,g:I=(-1,1)\to\overline{\mathfrak{R}}$ defined by f(x)=4f(x)=-4 and $g(x)=\pi x$. Then, the set $(f,g)_{I}$ is an open strong cell but not an open refined strong cell. Moreover, $(f,g)_{I}$ cannot be expressed as a finite union of open refined strong cells.
Lemma 2.11 ([5]).
Let $\mathcal{M}$ be a weakly o-minimal expansion of an ordered field. Then $\mathcal{M}$ is a real closed field.
Fact 2.12 ([9]).
Suppose that $\mathcal{M}=(M,<,+,\cdot,\ldots)$ is an o-minimal expansion of a real closed field. Let $U$ be a definable bounded open subset of $M^{n}$. Then, there exists a finite collection of open cells in $M^{n}$ whose union is $U$.
Theorem 2.13.
Suppose that $\mathcal{M}=(M,<,+,\cdot,\ldots)$ is a non-valuational weakly o-minimal expansion of a real closed field. Let $U$ be a definable bounded open subset of $M^{n}$. Then, there exists a finite collection of open strong cells in $M^{n}$ whose union is $U$.
Proof.
Let $\mathcal{C}$ be the smallest refined strong cell decomposition of $M^{n}$ partitioning $U$ such that $U=\bigcup_{i=1}^{k}C_{i}$ and $\overline{U}=\bigcup_{i=1}^{k}\overline{C_{i}}$ with $C_{1},\ldots,C_{k}\in\mathcal{C}$. By Remark 2.7, the set $\overline{U}$ is a definable bounded open subset of $\overline{M}^{n}$. Then, by Fact 2.12, there exist open cells $D_{1},\dots,D_{l}$ such that $\overline{U}=D_{1}\cup\dots\cup D_{l}$. Since each $D_{i}\cap M^{n}$ is an open strong cell and ${U}=(D_{1}\cap M^{n})\cup\dots\cup(D_{l}\cap M^{n})$, this completes the proof. ∎
3. Weakly o-minimal expansions of ordered groups
Let $\mathcal{M}$ be a non-vaulational weakly o-minimal expansion $\mathcal{M}=(M,<,+,0,\cdots)$ of an ordered group. We say that $\mathcal{M}$ is semisemi-boundedbounded if there exists no definable bijection between a bounded interval and a unbounded interval.
Theorem 3.1 ([3]).
Let $\mathcal{M}$ be a semi-bounded o-minimal expansion $\mathcal{M}=(M,<,+,0,\cdots)$ of an ordered group. Then every non-empty definable open set is a finite union of open cells.
Theorem 3.2.
Consider a semi-bounded non-valuational weakly o-minimal expansion $\mathcal{M}=(M,<,+,\dots)$ of an ordered group. Let $U$ be a definable open subset of $M^{n}$. Then there exist finitely many strong open cells whose union is $U$.
Proof.
Take the smallest refined strong cell decomposition $\mathcal{C}$ of $M^{n}$ partitioning $U$ such that $U=\bigcup_{i=1}^{k}C_{i}$ and $\overline{U}=\bigcup_{i=1}^{k}\overline{C_{i}}$ with $C_{1},\ldots,C_{k}\in\mathcal{C}$. By Remark 2.7, the set $\overline{U}$ is a definable open subset of $\overline{M}^{n}$. Then, by Theorem 3.1, there exist open cells $D_{1},\dots,D_{l}$ such that $\overline{U}=D_{1}\cup\dots\cup D_{l}$. Since each $D_{i}\cap M^{n}$ is an open strong cell and ${U}=(D_{1}\cap M^{n})\cup\dots\cup(D_{l}\cap M^{n})$, this completes the proof. ∎