### v2

# Open cell property in weakly o-minimal structures

Abstract.
Every bounded definable open set is a union of finitely many open strong cells in a weakly o-minimal expansion of a real closed field. We prove this fact and another theorem similar to it.
Key words and phrases:
Weakly o-minimal structures, open strong cells and open cell property. The first author was partially supported by JSPS KAKENHI Grant Number 23K03095.
2020 Mathematics Subject Classification:
Primary 03C64
1. Introduction
Throughout this paper, “definable” means “definable possibly with parameters” and we assume that a structure $\mathcal{M}=(M,<,\dots)$ is a dense linear order without endpoints.
A subset $A$ of $M$ is said to be convex if $a,b\in A$ and $c\in M$ with $a<c<b$ then $c\in A$. Moreover if $A=\emptyset$ or $\inf A,{\rm sup}A\in M\cup\{-\infty,+\infty\}$, then $A$ is called an interval in $M$. We say that $\mathcal{M}$ is o-minimal (weakly o-minimal) if every definable subset of $M$ is a finite union of intervals (convex sets), respectively. Weakly o-minimality was introduced by Dickmann in [3].
Wilkie ([12]) proved that in o-minimal expansions of a real closed field every bounded definable open set is a finite union of open cells. Andrews ([1]) proved that if an o-minimal structure admits CE-cell decomposition then any definable open set is expressed as a finite union of definable open cells. Edmundo et al ([6]) proved in semi-bounded o-minimal expansion of an ordered group, any definable open set is covered by a finite union of open cells. The relation between strong cell decomposition in weakly o-minimal structures and the open cell property in o-minimal structures is proved by [7]. We consider a generalization of the above results in weakly o-minimal structures.
Readers are expected to be familiar with fundamental results of o-minimality and weak o-minimality; see, for example, [2], [4], [9], [10].
2. Weakly o-minimal expansions of real closed fields
Let $\mathcal{M}=(M,<,\ldots)$ be a weakly o-minimal expansion of a dense linear order without endpoints. For any subsets $C,D$ of $M$, we write $C<D$ if $c<d$ whenever $c\in C$ and $d\in D$. A pair $\langle C,D\rangle$ of non-empty subsets of $M$ is called a cut in $\mathcal{M}$ if $C<D,C\cup D=M$ and $D$ has no lowest element. A cut $\langle C,D\rangle$ is said to be definable in $\mathcal{M}$ if the sets $C$ and $D$ are definable in $\mathcal{M}$. The set of all cuts definable in $\mathcal{M}$ will be denoted by $\overline{M}$. Note that we have $M=\overline{M}$ if $\mathcal{M}$ is o-minimal. We define a linear ordering on $\overline{M}$ by $\langle C_{1},D_{1}\rangle<\langle C_{2},D_{2}\rangle$ if and only if $C_{1}\subsetneq C_{2}$. Then we may treat $(M,<)$ as a substructure of $(\overline{M},<)$ by identifying an element $a\in M$ with the definable cut $\langle(-\infty,a],(a,+\infty)\rangle$. Moreover, if $\mathcal{M}=(M,<,+,\ldots)$ is an expansion of an ordered group $(M,<,+)$, then the cut $\langle C,D\rangle$ is said to be valuational if there exists $\varepsilon>0$ such that for all $x\in C$ and $y\in D$, we have $y-x>\varepsilon$. Otherwise the cut $\langle C,D\rangle$ is said to be non-valuational. The structure $\mathcal{M}$ is said to be non-valuational if all definable cuts in $\mathcal{M}$ are non-valuational.
We equip $M$ ($\overline{M}$) with the interval topology (the open intervals form a base), and each product $M^{n}$ ($(\overline{M})^{n}$) with the corresponding product topology, respectively.
Recall the notion of definable functions from [11]. Let $n$ be a positive integer and $A\subseteq M^{n}$ definable. A function $f:A\to\overline{M}$ is said to be definable if the set $\{\langle x,y\rangle\in M^{n+1}:x\in A,y<f(x)\}$ is definable. A function $f:A\to\overline{M}\cup\{-\infty,+\infty\}$ is said to be definable if $f$ is a definable function from $A$ to $\overline{M}$, $f(x)=$-$\infty$ for all $x\in A$, or $f(x)=+\infty$ for all $x\in A$.
We recall the notion of strong cells from [10].
Definition 2.1.
For every $m\in\mathbb{N}_{+}$, we define, by induction, strong cells $C\subseteq M^{m}$ and their completions $\overline{C}\subseteq\overline{M}^{m}$.
  1. (1)
    Any singleton of $M$ is a 0-strong cell in $M$ and is equal to its completion.
  2. (2)
    Any non-empty open convex definable subset of $M$ is a 1-strong cell in $M$. If $C\subseteq M$ is a 1-strong cell, then we define its completion by $\overline{C}:=\{x\in\overline{M}:(\exists a,b\in C)(a<x<b)\}$.
    Let $m\in\mathbb{N}_{+}$, $k\leq m$ and suppose that we have already defined $k$-strong cells in $M^{m}$ and their completions in $\overline{M}^{m}$.
  3. (3)
    If $C\subseteq M^{m}$ is a $k$-strong cell and $f:C\to M$ is a continuous definable function which has a unique continuous extension $\overline{f}:\overline{C}\to\overline{M}$, then the graph of $f$ which is denoted by $\Gamma(f)_{C}$, is a $k$-strong cell in $M^{m+1}$ and its completion in $\overline{M}^{m+1}$ is defined as $\Gamma(\overline{f})_{\overline{C}}$.
  4. (4)
    If $C\subseteq M^{m}$ is a $k$-strong cell and $f,g:C\to\overline{M}\cup\{\pm\infty\}$ is continuous definable functions such that $f,g$ have continuous extensions $\overline{f},\overline{g}:\overline{C}\to\overline{M}\cup\{\pm\infty\}$, where $\overline{f}(\overline{a})<\overline{g}(\overline{a})$ for any $\overline{a}\in\overline{C}$, then the set $(f,g)_{C}:=\{\langle\overline{a},b\rangle\in C\times M:f(\overline{a})<b<g(\overline{a})\}$ is called a $(k+1)$-strong cell in $M^{m+1}$. The completion of $(f,g)_{C}$ in $\overline{M}^{m+1}$ is defined as $(\overline{f},\overline{g})_{\overline{C}}:=\{\langle\overline{a},b\rangle\in\overline{C}\times\overline{M}:\overline{f}(\overline{a})<b<\overline{g}(\overline{a})\}$.
  5. (5)
    Every $k$-strong cell is of the form given in (1) through (4). We say that $C\subseteq M^{m}$ is a strong cell in $M^{m}$ if there exists a non-negative integer $k$ such that $C$ is a $k$-strong cell in $M^{m}$.
A function $f:C\to\overline{M}$ is said to be a cell-defining function if $\Gamma(f)_{C}$ is a strong cell or there is a definable function $g:C\to\overline{M}$ so that $(f,g)_{C}$ or $(g,h)_{C}$ is a strong cell.
A strong cell $C\subseteq M^{m}$, $m\geq 2$, is called a refined strong cell if each of the cell-defining functions appearing in its definition assumes all its values in one of the following sets: $M$, $\overline{M}\setminus M$, $\{-\infty\}$, or $\{+\infty\}$. Refined strong cells in $M$ coincide with strong cells in $M$. Note that the definitions of strong cells in [10] and [11] are not identical. A strong cell in [11] is called a refined strong cell in [10]. In this paper, we follow the terminology of [10].
Let $C\subseteq M^{m}$ be a strong cell and $f:C\to\overline{M}$ definable. The function $f$ is said to be strongly continuous if it has a unique continuous extension $\overline{f}:\overline{C}\to\overline{M}$.
Definition 2.2.
Any finite partition of $M$ into singletons and convex open sets definable in $\mathcal{M}$ is called a strong cell decomposition of $M$. A finite partition $\mathcal{C}$ of $M^{m+1}$ into strong cells is said to be a strong cell decomposition of $M^{m+1}$ if $\pi[\mathcal{C}]:=\{\pi[C]:C\in\mathcal{C}\}$ is a strong cell decomposition of $M^{m}$. Here $\pi:M^{m+1}\to M^{m}$ denotes the projection dropping the last coordinate. A strong cell decomposition $\mathcal{C}$ of $M^{m}$ partitions $X\subseteq M^{m}$ if for any $C\in\mathcal{C}$, we have $C\subseteq X$ or $C\cap X=\emptyset$. The structure $\mathcal{M}$ has the strong cell decomposition property if for any positive integers $m$, $k$ and any definable sets $X_{1},\ldots,X_{k}\subseteq M^{m}$, there exists a decomposition of $M^{m}$ into finitely many strong cells partitioning each of the sets $X_{1},\ldots,X_{k}$. Similarly, we can define the refined strong cell decomposition property.
The following result follows from Lemma 2.6, Corollary 2.16, and Corollary 2.17 of [10].
Fact 2.3.
Let $\mathcal{M}=(M,<,+,\ldots)$ be a weakly o-minimal expansion of an ordered group. Then the structure $\mathcal{M}$ is non-valuational if and only if it has the refined strong cell decomposition property.
The following definition appears in [11].
Definition 2.4.
Let $\mathcal{M}=(M,<,\ldots)$ be a weakly o-minimal structure with the refined strong cell decomposition property. For any $m\in\mathbb{N}_{+}$ and $i_{1},\dots,i_{m}\in\{0,1\}$, we define basic $\langle i_{1},\dots,i_{m}\rangle$-cells in $\overline{M}^{m}$ inductively.
If $1\leq j_{1}<\dots<j_{k}\leq m$, let
$\rho^{m}_{j_{1},\dots,j_{k}}:\overline{M}^{m}\to\overline{M}^{k}$
denote the projection onto the coordinates $j_{1},\dots,j_{k}$. When $m$ is clear from the context, we simply write $\rho_{j_{1},\dots,j_{k}}$.
  1. (1)
    A singleton of $\overline{M}^{m}$ is called a basic $\langle 0,\dots,0\rangle$-cell in $\overline{M}^{m}$.
  2. (2)
    If $C\subseteq M$ is a strong $\langle 1\rangle$-cell, then $\overline{C}$ is called a basic $\langle 1\rangle$-cell. Note that
    $\rho_{1}(\overline{C})\cap M=\overline{C}\cap M=C$
    is an open strong cell in $M$.
  3. (3)
    If $C=\{\overline{a}\}\subseteq\overline{M}^{m}$ and $I$ is a basic $\langle 1\rangle$-cell in $\overline{M}$, then $C\times I$ is called a basic $\langle 0,\dots,0,1\rangle$-cell in $\overline{M}^{m+1}$. Note that
    $\rho_{m+1}(C\times I)\cap M=I\cap M$
    is an open strong cell in $M$.
  4. Assume that $i_{1},\dots,i_{m}\in\{0,1\}$ with $i_{1}+\dots+i_{m}>0$, and suppose that basic $\langle i_{1},\dots,i_{m}\rangle$-cells in $\overline{M}^{m}$ have already been defined. Let
    $\{j_{1},\dots,j_{k}\}=\{j\in\{1,\dots,m\}:i_{j}=1\},\qquad j_{1}<\dots<j_{k},$
    and suppose that for every basic $\langle i_{1},\dots,i_{m}\rangle$-cell $C\subseteq\overline{M}^{m}$, the set
    $\rho_{j_{1},\dots,j_{k}}(C)\cap M^{k}$
    is an open strong cell in $M^{k}$.
  5. (4)
    Let $C\subseteq\overline{M}^{m}$ be a basic $\langle i_{1},\dots,i_{m}\rangle$-cell, and set
    $D:=\rho_{j_{1},\dots,j_{k}}(C)\cap M^{k},$
    which is an open strong cell in $M^{k}$. If $f$ is a strongly continuous definable function from $D$ to $M$ or from $D$ to $\overline{M}\setminus M$, then
    $\Gamma\bigl(\overline{f}\circ\rho_{j_{1},\dots,j_{k}}\!\restriction_{C}\bigr)$
    is called a basic $\langle i_{1},\dots,i_{m},0\rangle$-cell in $\overline{M}^{m+1}$. Note that
    $\rho_{j_{1},\dots,j_{k}}\!\left(\Gamma\bigl(\overline{f}\circ\rho_{j_{1},\dots,j_{k}}\!\restriction_{C}\bigr)\right)\cap M^{k}=D$
    is an open strong cell in $M^{k}$.
  6. (5)
    Let $C\subseteq\overline{M}^{m}$ be a basic $\langle i_{1},\dots,i_{m}\rangle$-cell, and set
    $D:=\rho_{j_{1},\dots,j_{k}}(C)\cap M^{k},$
    an open strong cell in $M^{k}$. If $f,g:D\to\overline{M}\cup\{\pm\infty\}$ are strongly continuous definable functions such that all values of $f$ and $g$ lie in one of the sets
    $\{-\infty\},\quad M,\quad\overline{M}\setminus M,\quad\{+\infty\},$
    and
    $\overline{f}(\overline{x})<\overline{g}(\overline{x})\quad\text{for all }\overline{x}\in\overline{D},$
    then the set
    $(\overline{f}\circ\rho_{j_{1},\dots,j_{k}},\overline{g}\circ\rho_{j_{1},\dots,j_{k}})_{C}:=\left\{\langle\overline{a},b\rangle\in C\times\overline{M}:(\overline{f}\circ\rho_{j_{1},\dots,j_{k}})(\overline{a})<b<(\overline{g}\circ\rho_{j_{1},\dots,j_{k}})(\overline{a})\right\}$
    is called a basic $\langle i_{1},\dots,i_{m},1\rangle$-cell in $\overline{M}^{m+1}$. Note that
    $\rho_{j_{1},\dots,j_{k}}\left[(\overline{f}\circ\rho_{j_{1},\dots,j_{k}},\overline{g}\circ\rho_{j_{1},\dots,j_{k}})_{C}\right]\cap M^{k+1}=(f,g)_{D}$
    is an open strong cell in $M^{k+1}$.
In the standard way, we define a basic cell decomposition of $\overline{M}^{m}$ as a finite partition of $\overline{M}^{m}$ into basic cells.
For any $n\in\mathbb{N}_{+}$ and any refined strong cell $C\subseteq M^{n}$, we denote by $R_{C}$ an $n$-ary relational symbol. We interpret $R_{C}$ in $\overline{M}^{n}$ as $\overline{C}$, the completion of $C$. According to Section 2 of [11], the structure $\overline{\mathcal{M}}:=(\overline{M},<,(R_{C}:C\text{ is a refined strong cell}))$ is o-minimal and is called the canonical o-minimal extension of $\mathcal{M}$. If $X\subseteq\overline{M}^{m}$ is a set definable in $\overline{\mathcal{M}}$, then $X\cap M^{m}$ is definable in $\mathcal{M}$. If additionally $Y\subseteq M^{m}$ is definable in $\mathcal{M}$, then $X\cap Y$ is definable in $\mathcal{M}$.
The following result follows from Proposition 2.3 of [11].
Fact 2.5.
Assume that $\mathcal{M}=(M,<,\dots)$ is a weakly o-minimal structure with the strong cell decomposition property and $\overline{\mathcal{M}}=(\overline{M},<,\dots)$ is its canonical o-minimal extension. Let $m,k\in\mathbb{N}_{+}$. If $X_{1},\dots,X_{k}\subseteq\overline{M}^{m}$ are sets definable in $\overline{{\mathcal{M}}}$, then there exists a basic cell decomposition of $\overline{M}^{m}$ into finitely many basic cells partitioning each of the sets $X_{1},\ldots,X_{k}$.
Proposition 2.6.
Suppose that $\mathcal{M}=(M,<,\ldots)$ is a weakly o-minimal structure with the refined strong cell decomposition property. Let $n\in\mathbb{N}_{+}$. Suppose that $X$ is a definable open subset of $M^{n}$. Then there exists an $\overline{{\mathcal{M}}}$-definable open set $Y\subseteq\overline{M}^{n}$ such that $X=Y\cap M^{n}$.
Proof.
Take a refined strong cell decomposition of $X$, so that
$X=C_{1}\cup\dots\cup C_{k}.$
By Fact 2.5, there exists a basic cell decomposition $\mathcal{C}$ of $\overline{M}^{n}$ partitioning each of the sets $\overline{C_{1}},\dots,\overline{C_{k}}$. Define
$\hat{X}=\bigcup\left\{D\in\mathcal{C}:\exists i\ (D\subseteq\overline{C_{i}})\right\}\;\cup\;\bigcup\left\{D\in\mathcal{C}:\bigl(\forall i\ (D\cap\overline{C_{i}}=\emptyset)\bigr)\wedge(D\cap M^{n}=\emptyset)\right\}.$
[…]
By the definition of $\hat{X}$, the inclusion $X\supseteq Y\cap M^{n}$ is clear.
[…]
If $C\subseteq\overline{C_{i}}$ for some $i$, then $B\subseteq\overline{C_{i}}\subseteq\hat{X}$, a contradiction. Thus for every $i$ we have $C\cap\overline{C_{i}}=\emptyset$, and hence $C\cap X=\emptyset$.
First, consider the case where $C$ is a basic $\langle 0,\dots,0,1\rangle$-cell. Then there exists an open interval $I\subseteq\overline{M}$ such that
$C=(a_{1},\dots,a_{n-1})\times I.$
Since $a\in X\subseteq M^{n}$ and $a\in\operatorname{cl}_{\overline{M}}(C)$, every open box $U$ in $M^{n}$ containing $a$ satisfies $C\cap U\neq\emptyset$, equivalently $(C\cap M^{n})\cap U\neq\emptyset$. Because $C\cap X=\emptyset$, we have $(C\cap M^{n})\cap X=\emptyset$. Hence $a\notin\operatorname{int}_{M}(X)$, contradicting that $X$ is open in $M^{n}$.
Next, consider the case where $C$ is a basic $\langle i_{1},\dots,i_{n-1},0\rangle$-cell and $i_{1}+\dots+i_{n-1}>0$. That is, there exists a basic $\langle i_{1},\dots,i_{n-1}\rangle$-cell $\hat{D}\subseteq\overline{M}^{n-1}$, and if we set
$\{j_{1},\ldots,j_{k}\}=\{\,j\in\{1,\dots,n-1\}:i_{j}=1\,\},\qquad j_{1}<\dots<j_{k},$
then
$D=\rho_{j_{1},\ldots,j_{k}}(\hat{D})\cap M^{k}$
is an open strong cell in $M^{k}$. Moreover, there exists a strongly continuous definable function $f$ from $D$ to $M$ or from $D$ to $\overline{M}\setminus M$ such that
[…]
Assume that $f:D\to M$. Recall that $C\cap\overline{C_{i}}=\emptyset$ for every $1\leq i\leq k$. If $C\cap M^{n}=\emptyset$, $C\subseteq\hat{X}$ by the definition of $\hat{X}$. Hence $B\subseteq C\subseteq\hat{X}$, a contradiction.
Assume that $C\cap M^{n}\neq\emptyset$. In the construction of a function-type cell in the basic cells, all cell-defining functions take values in $M$. We claim that $a\in\operatorname{cl}_{\overline{M}}(C\cap M^{n})$. Let $U\subseteq M^{n}$ be an arbitrary open box containing $a$. Since $a\in\operatorname{cl}_{\overline{M}}(C)$, there exists a point $x=(x_{1},\dots,x_{n})\in\overline{U}\cap C$. Set $\tilde{C}=\rho_{j_{1},\ldots,j_{k}}(C)$ and $V=\rho_{j_{1},\ldots,j_{k}}(U)$. Then $\tilde{C}$ is open in $\overline{M}^{k}$, and $V$ is an open box in $M^{k}$. Let $\tilde{x}=(x_{j_{1}},\dots,x_{j_{k}})$. Then $\tilde{x}\in\overline{V}\cap\tilde{C}$. Since $\tilde{C}$ is open in $\overline{M}^{k}$, there exists an open box $W$ in $M^{k}$ such that $\tilde{x}\in\overline{W}\subseteq\overline{V}\cap\tilde{C}$. Choose $\tilde{y}\in W$. Since $f:D\to M$, there exists $y\in(C\cap M^{n})\cap U$ such that $\rho_{j_{1},\ldots,j_{k}}(y)=\tilde{y}$. Therefore $(C\cap M^{n})\cap U\neq\emptyset$, and hence $a\in\operatorname{cl}_{\overline{M}}(C\cap M^{n})$. However, since $C\cap X=\emptyset$, we have $(C\cap M^{n})\cap X=\emptyset$, and hence $a\notin\operatorname{int}_{M}(X)$. This contradicts the assumption that $X$ is open in $M^{n}$.
Assume instead that $f:D\to\overline{M}\setminus M$. Then $C\cap M^{n}=\emptyset$, and similarly $C\subseteq\hat{X}$. Hence $B\subseteq C\subseteq\hat{X}$, a contradiction.
Therefore the function-type case leads to a contradiction.
Finally, consider the case where $C$ is a basic $\langle i_{1},\dots,i_{n-1},1\rangle$-cell and $i_{1}+\dots+i_{n-1}>0$. As before, there exist a basic $\langle i_{1},\dots,i_{n-1}\rangle$-cell $\hat{D}\subseteq\overline{M}^{n-1}$ and an open strong cell $D\subseteq M^{k}$. Moreover, there exist strongly continuous functions $f,g:D\to\overline{M}\cup\{\pm\infty\}$ such that all values of $f$ and $g$ lie in exactly one of $\{-\infty\},\ M,\ \overline{M}\setminus M,\ \{\infty\}$, and $\overline{f}(x)<\overline{g}(x)$ for any $x\in\overline{D}$, and
$C=(\overline{f}\circ\rho_{j_{1},\ldots,j_{k}},\ \overline{g}\circ\rho_{j_{1},\ldots,j_{k}})_{\hat{D}}$.
Recall that $C\cap\overline{C_{i}}=\emptyset$ for every $1\leq i\leq k$. If $C\cap M^{n}=\emptyset$, $C\subseteq\hat{X}$ by the definition of $\hat{X}$. Hence $B\subseteq C\subseteq\hat{X}$, a contradiction.
Assume $C\cap M^{n}\neq\emptyset$. As in the previous case, we have $a\in\operatorname{cl}_{\overline{M}}(C\cap M^{n})$. Hence every open box $U$ in $M^{n}$ containing $a$ meets $C\cap M^{n}$. Because $C\cap X=\emptyset$, we have $(C\cap M^{n})\cap X=\emptyset$, so $a\notin\operatorname{int}_{M}(X)$, contradicting openness of $X$ in $M^{n}$. This completes the proof. ∎
We recall the notion of open cell property from [1]. An o-minimal structure $\mathcal{M}=(M,<,\dots)$ is said to have the open cell property if every non-empty definable open subset of $M^{n}$ is a union of finitely many open cells. A weakly o-minimal structure $\mathcal{M}=(M,<,\dots)$ is said to have the open cell property if every non-empty definable open subset of $M^{n}$ is a union of finitely many open strong cells.
The following example shows that the open cell property fails for refined strong cells.
Example 2.7.
Let $R_{\mathrm{alg}}$ denote the ordered field of real algebraic numbers and let the unary predicate symbol $P$ be interpreted by the convex set $S=(-\pi,\pi)\cap R_{\mathrm{alg}}$. Then, by Proposition 2.1 of [9], the structure $\mathfrak{R}=(R_{\mathrm{alg}},<,+,\cdot,P)$ is non-valuational weakly o-minimal. So the structure $\mathfrak{R}$ has the refined strong cell decomposition property. Consider the functions $f,g:I=(-1,1)\to\overline{R}_{\mathrm{alg}}$ defined by f(x)=4f(x)=-4 and $g(x)=\pi x$. The set $(f,g)_{I}$ is an open strong cell but not a refined strong cell because $g(0)=0\in R_{\mathrm{alg}}$ and $g(x)\notin R_{\mathrm{alg}}$ for $x\neq 0$. Moreover, $(f,g)_{I}$ cannot be expressed as a finite union of open refined strong cells.
Fact 2.8 ([12, Theorem 1.3]).
Suppose that $\mathcal{M}=(M,<,+,\cdot,\ldots)$ is an o-minimal expansion of a real closed field. Let $U$ be a definable bounded open subset of $M^{n}$. Then, there exists a finite collection of open cells in $M^{n}$ whose union is $U$.
Theorem 2.9.
Suppose that $\mathcal{M}=(M,<,+,\cdot,\ldots)$ is a non-valuational weakly o-minimal expansion of a real closed field. Let $U$ be a definable bounded open subset of $M^{n}$. Then, there exists a finite collection of open strong cells in $M^{n}$ whose union is $U$.
Proof.
By Fact 2.3, the structure $\mathcal{M}$ has the refined strong cell decomposition property. Since $U$ is a definable bounded open subset of $M^{n}$, by Proposition 2.6, there exists an $\overline{{\mathcal{M}}}$-definable bounded open set $V\subseteq\overline{M}^{n}$ such that $U=V\cap M^{n}$. By Fact 2.8, there exist open cells $D_{1},\dots,D_{l}$ such that $V=D_{1}\cup\dots\cup D_{l}$. Because each $D_{i}\cap M^{n}$ is an open strong cell and ${U}=(D_{1}\cap M^{n})\cup\dots\cup(D_{l}\cap M^{n})$, this completes the proof. ∎
3. Weakly o-minimal expansions of ordered groups
Let $\mathcal{M}=(M,<,+,\cdots)$ be a non-valuational weakly o-minimal expansion of an ordered group. If $\mathcal{M}$} is o-minimal, we say that $\mathcal{M}$ is semi-bounded if there exists no definable bijection between a bounded interval and an unbounded interval [5].
Theorem 3.1 ([6]).
Let $\mathcal{M}=(M,<,+,\cdots)$ be a semi-bounded o-minimal expansion of an ordered group. Then every non-empty definable open set is a finite union of open cells.
Theorem 3.2.
Consider a non-valuational weakly o-minimal expansion $\mathcal{M}=(M,<,+,\dots)$ of an ordered group. Let $U$ be a definable open subset of $M^{n}$ and assume that $\overline{{\mathcal{M}}}$ is semi-bounded. Then there exist finitely many open strong cells whose union is $U$.
Proof.
By Proposition 2.6, there exists an $\overline{{\mathcal{M}}}$-definable open subset $V$ of $\overline{M}^{n}$ such that $U=V\cap M^{n}$. By Theorem 3.1, there exist open cells $D_{1},\dots,D_{l}$ such that $V=D_{1}\cup\dots\cup D_{l}$. Since each $D_{i}\cap M^{n}$ is an open strong cell and ${U}=(D_{1}\cap M^{n})\cup\dots\cup(D_{l}\cap M^{n})$, this completes the proof. ∎
We obtained an improvement of Theorem 3.2 inspired by a private communication with Fujita [8].
Theorem 3.3.
Let $\mathcal{M}=(M,<,+,\dots)$ be a non-valuational weakly o-minimal expansion of an ordered group. Then the following three conditions are equivalent.
(1) For any bounded definable set $I\subseteq M$ and for every $\mathcal{M}$-definable function $f:I\to\overline{M}$, $f(I)$ is bounded.
(2) The canonical o-minimal extension $\overline{\mathcal{M}}$ of $\mathcal{M}$ is semi-bounded.
(3) Any $\mathcal{M}$-definable open set is a finite union of open strong cells.
Proof.
$(1)\Rightarrow(2)$ Assume that (2) does not hold. Then we have a bounded open interval $J_{1}$, an unbounded interval $J_{2}$ and an $\overline{{\mathcal{M}}}$-definable bijection $g:J_{1}\to J_{2}$. Since $\overline{{\mathcal{M}}}$ is o-minimal and by the monotonicity theorem, we may assume $g$ is continuous and strictly monotone. Set $I=J_{1}\cap M$, $f=g|I$. Then $f$ is $\mathcal{M}$-definable. Since $M$ is dense in $\overline{M}$, $f(I)$ is unbounded. This contradicts (1).
$(2)\Rightarrow(1)$ We show the contradiction. Suppose that $I$ is a bounded subset of $M$ and $f:I\to\overline{M}$ is a function definable in $\mathcal{M}$. Assume that $f(I)$ is unbounded. By Proposition 3.1 of [11], we can decompose $I$ into $I=X\cup J_{1}\cup\dots\cup J_{k}$ such that $X$ is finite, each $J_{i}$ is open convex set and $f|J_{i}$ is monotone and strongly continuous. We have the fact that some $f(J_{i})$ is unbounded. Replacing $f$ by $f|J_{i}$, we can assume that $f$ is monotone and strongly continuous. Then the unique extension $\overline{f}:\overline{I}\to\overline{M}$ is monotone and continuous. This is an $\overline{{\mathcal{M}}}$-definable bijection between a bounded open interval $\overline{I}$ and an unbounded open interval $\overline{f}(\overline{I})$. This is a contradiction.
$(2)\Rightarrow(3)$ By Theorem 3.2, this implication follows.
$(3)\Rightarrow(2)$ Assume that (2) does not hold. Then we have an $\overline{{\mathcal{M}}}$-definable bijection $f:J_{1}\to J_{2}$ between a bounded open interval $J_{1}$ and an unbounded interval $J_{2}$. Since $\overline{{\mathcal{M}}}$ is o-minimal and by the monotonicity theorem, we may assume that $f$ is strictly monotone and continuous. We may further assume that $J_{2}$ is bounded below and $f$ is strictly decreasing. By shifting, we can set $J_{1}=(0,b)$. Let $Y$ be the $\overline{{\mathcal{M}}}$-definable open set
$Y=\{(x,y)\in\overline{M}^{2}:(-b<x<b)\land((x<0)\to(y<f(-x)))\land((x>0)\to(y<f(x)))\}$.
Set $X=Y\cap M^{n}$. Then $X$ is $\mathcal{M}$-definable. We prove that $X$ is not a finite union of open strong cells.
Assume $X$ is a finite union of open strong cells. Note that $\{0\}\times M$ is contained in $X$. Thus there exists an open strong cell $C$ and $r\in M$ such that $\{0\}\times(r,\infty)\subseteq C\subseteq X$. Since $\{0\}\times(r,\infty)\subseteq C$, we can find a definable open convex set $I$ with $0\in I$ and a strongly continuous definable function $g:I\to\overline{M}\cup\{-\infty\}$ such that $C=\{(x,y)\in M^{2}:x\in I,y>g(x)\}$. Hence $C$ contains a point $(x_{0},y_{0})$ with $0<x_{0}<b$ and $y_{0}>f(x_{0})$. It contradicts $C\subseteq X$. ∎